OpenStudy (anonymous):

A student studying for a vocabulary test knows the meanings of 16 words from a list of 20 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows?

OpenStudy (anonymous):

at least 8 means 8 or 9 or 10 so you have to compute these separately and add

OpenStudy (anonymous):

But how do I do that?

OpenStudy (anonymous):

number of ways to pick 10 out of 20 is \(\binom{20}{10}=184756\)so that is your denominator

OpenStudy (anonymous):

so the numerator would be (8 over 10)?? But what about the 16?

OpenStudy (anonymous):

lets do 10 first, because that is the easiest. number of ways to choose 10 out of the 16 that he knows is \(\binom{16}{10}=8008\) so that is your numerator for the probability that he knows all ten

OpenStudy (anonymous):

i.e. the probability that he knows all ten is \[\frac{\dbinom{16}{10}}{\dbinom{20}{10}}=\frac{8008}{184765}\]

OpenStudy (anonymous):

now for nine out of the ten. that means he has gotten 9 form the 16 he knows and 1 from the other 4 he does not, so that answer would be \[\frac{\dbinom{16}{9}\times \dbinom{4}{1}}{\dbinom{20}{10}}\]

OpenStudy (anonymous):

and finally for 8 out of ten, he got 8 from the 16, 2 from the other 4, so that would be \[\frac{\dbinom{16}{8}\times \dbinom{4}{2}}{\dbinom{20}{10}}\] i will leave it to you to compute these numbers and add

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

yw